# Understanding and Improving Shampoo and SOAP via Kullback–Leibler Minimization

### v2

@@ 4 Efficient Implementation: KL-Shampoo with QR Decomposition
$\smash{\min_{{\bm{\lambda}}_{a},{\bm{\lambda}}_{b}}\mathrm{KL}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big)}$ with preconditioner $\smash{{\bm{S}}\!=\!({\bm{Q}}_{a}\mathrm{Diag}({\bm{\lambda}}_{a}){\bm{Q}}_{a}^{\top})\!\otimes\!({\bm{Q}}_{b}\mathrm{Diag}({\bm{\lambda}}_{b}){\bm{Q}}_{b}^{\top})}$ should satisfy the following condition.
@@ 5 Interpreting and Improving SOAP via KL Minimization
(SOAP and KL-SOAP’s covariance estimation for augmented eigenvalues $\bm{d}$}) The optimal solution of KL minimization: $\min_{{\bm{d}}}\mathrm{KL}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big)$ with preconditioner ${\bm{S}}={\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$ is ${\bm{d}}^{*}=\mathbb{E}\left[\left(\mathrm{vec}({\bm{Q}}_{a}^{\top}{\bm{G}}{\bm{Q}}_{b})\right)^{\odot 2}\right]=\mathbb{E}\left[\hat{{\bm{g}}}^{\odot 2}\right]$, where ${\bm{d}}\in\mathcal{R}^{d_{a}d_{b}\times 1}$ is viewed as an augmented eigenvalue vector, $\hat{{\bm{g}}}={\bm{Q}}^{\top}{\bm{g}}$ is defined at the update of SOAP (see Eq.˜2), and ${\bm{Q}}={\bm{Q}}_{a}\otimes{\bm{Q}}_{b}$ can be an outdated eigenbasis of (KL-)Shampoo’s preconditioner.
@@ 6 Experimental Setup and Empirical Evaluations
Figure 5: Empirical results—based on random search with 150 runs per method—demonstrate the advantages of KL-Shampoo (two-sided) over other Shampoo variants under comparable settings, including Shampoo with $p=\nicefrac{{1}}{{2}}$ (one-sided, no grafting), F-Shampoo (two-sided, Frobenius-norm–based), and VN-Shampoo (trace scaling, two-sided von-Neumann-divergence-based). We make these variants practical by incorporating a QR step and an EMA scheme for eigenvalue estimation (Fig.˜3). To ensure a fair comparison and minimize implementation bias, we implement Shampoo, F-Shampoo, and VN-Shampoo ourselves, aligning them closely with KL-Shampoo. See Fig.˜10 (Appx.˜H) for a detailed comparison between KL-Shampoo and VN-Shampoo.
@@ Appendix E Proof of ˜5
Appendix F two-sided Shampoo Update based on Frobenius norm
@@
(Shampoo’s estimation scheme based on Frobenius norm) The optimal solution of Frobenius norm minimization $\min_{{\bm{S}}_{a},{\bm{S}}_{b}}\mathrm{Frob}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big):=\|\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]-{\bm{S}}\|_{\text{Frob}}$ with a two-sided precontioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}$ should satisfy the following condition.
@@
Remark 1: Although the solution can be obtained via singular value decomposition (SVD) (Van Loan & Pitsianis, 1993) on this outer product, $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$, it is computationally infeasible to compute the solution due to the high dimensionality of the product.
Remark 2: Notably, the optimal solution does not admit an analytical expression in the tensor-valued cases because there is no closed-form tensor extension of SVD.
@@
$=$ $\mathrm{Tr}\big({\bm{S}}_{a}^{2}\otimes{\bm{S}}_{b}^{2}-2\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(${\bm{S}}_{k}$ is symmetric for $k\in\{a,b\}$)}}$
$=$ $\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a Kronecker product)}}$
$=$ $\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big(({\bm{g}}{\bm{g}}^{\top})({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of the expectation)}}$
$=$ $\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}^{\top}\mathrm{vec}({\bm{S}}_{a}{\bm{G}}{\bm{S}}_{b})\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a Kronecker product)}}$
$=$ $\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{S}}_{a}{\bm{G}}{\bm{S}}_{b}\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a trace)}}$
@@ Appendix G Key Distinction between Shampoo with trace scaling and KL-Shampoo
(Shampoo’s estimation scheme with trace scaling) The optimal solution of von Neumann (VN) divergence (quantum relative entropy) minimization $\min_{{\bm{S}}_{a},{\bm{S}}_{b}}\mathrm{VN}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big):=\mathrm{Tr}({\bm{S}})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\mathrm{LogM}({\bm{S}})\big)+\text{const.}$ with a two-sided precontioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}$ should satisfy the following condition.
@@
${\bm{S}}_{a}^{*}$ $=\mathrm{Diag}\big(\mathbb{E}[{\bm{G}}{\bm{G}}^{\top}]\big)=\mathrm{Diag}\big(\mathbb{E}[{\bm{G}}^{\odot 2}\mathbf{1}]\big)$
${\bm{S}}_{b}^{*}$ $=\mathrm{Diag}\left(\frac{\mathbb{E}[{\bm{G}}^{\top}{\bm{G}}]}{\mathrm{Tr}\big(\mathbb{E}[{\bm{G}}{\bm{G}}^{\top}]\big)}\right)=\frac{\mathrm{Diag}\big(\mathbb{E}[\mathbf{1}^{\top}{\bm{G}}^{\odot 2}]\big)}{\mathrm{Tr}\big(\mathbb{E}[\mathbf{1}^{\top}{\bm{G}}^{\odot 2}\mathbf{1}]\big)}=\frac{\mathrm{Diag}\big(\mathbb{E}[\mathbf{1}^{\top}{\bm{G}}^{\odot 2}]\big)}{\sqrt{\mathrm{Tr}\big(\mathbb{E}[\mathbf{1}^{\top}{\bm{G}}^{\odot 2}\big)\mathrm{Tr}\big(\mathbb{E}[{\bm{G}}^{\odot 2}\mathbf{1}]\big)}}$
@@
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\mathrm{LogM}({\bm{S}})\big)+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}+{\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}\big)+\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big({\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}{\bm{g}}^{\top}\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}{\bm{g}}^{\top}\big({\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)\big]+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a}){\bm{G}}{\bm{I}}_{b}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{I}}_{a}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a}){\bm{G}}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$\quad=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}{\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a})\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$\quad=\mathrm{Tr}(\mathrm{ExpM}({\bm{P}}_{a}))\,\mathrm{Tr}(\mathrm{ExpM}({\bm{P}}_{b}))-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}{\bm{G}}^{\top}{\bm{P}}_{a}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}{\bm{P}}_{b}\big)\big]+\text{const.}$ (15)

### v3

@@ 4 Efficient Implementation: KL-Shampoo with QR Decomposition
$\smash{\min_{{\bm{\lambda}}_{a},{\bm{\lambda}}_{b}}\mathrm{KL}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big)}$ with preconditioner $\smash{{\bm{S}}\!=\!({\bm{Q}}_{a}\mathrm{Diag}({\bm{\lambda}}_{a}){\bm{Q}}_{a}^{\top})\!\otimes\!({\bm{Q}}_{b}\mathrm{Diag}({\bm{\lambda}}_{b}){\bm{Q}}_{b}^{\top})}$ satisfies this condition:
@@ 5 Interpreting and Improving SOAP via KL Minimization
(SOAP and KL-SOAP’s covariance estimation for augmented eigenvalues $\bm{d}$}) The optimal solution of KL minimization: $\min_{{\bm{d}}}\mathrm{KL}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big)$ with preconditioner ${\bm{S}}={\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$ is ${\bm{d}}^{*}=\mathbb{E}\left[\left(\mathrm{vec}({\bm{Q}}_{a}^{\top}{\bm{G}}{\bm{Q}}_{b})\right)^{\odot 2}\right]=\mathbb{E}\left[\hat{{\bm{g}}}^{\odot 2}\right]$, where ${\bm{d}}\in\mathcal{R}^{d_{a}d_{b}\times 1}$ is viewed as an augmented eigenvalue vector, $\hat{{\bm{g}}}={\bm{Q}}^{\top}{\bm{g}}$ is defined at the update of (KL-)SOAP (see Eq.˜2), and ${\bm{Q}}={\bm{Q}}_{a}\otimes{\bm{Q}}_{b}$ can be an outdated eigenbasis of (KL-)Shampoo’s preconditioner.
@@ 6 Experimental Setup and Empirical Evaluations
Figure 5: Empirical results—based on random search with 150 runs per method—demonstrate the advantages of KL-Shampoo’s (two-sided) estimation over other Shampoo variants under comparable settings for NN training, including Shampoo with $p=\nicefrac{{1}}{{2}}$ (one-sided, no grafting), F-Shampoo (two-sided, Frobenius-norm–based), and VN-Shampoo (trace scaling, two-sided von-Neumann-divergence-based). We make these variants practical by incorporating a QR step and an EMA scheme for eigenvalue estimation (Fig.˜3). To ensure a fair comparison and minimize implementation bias, we implement Shampoo, F-Shampoo, and VN-Shampoo ourselves, aligning them closely with KL-Shampoo. See Fig.˜10 (Appx.˜H) for a detailed comparison between KL-Shampoo and VN-Shampoo.
@@ Appendix E Proof of ˜5
Appendix F Two-sided Shampoo Scheme based on Frobenius norm
@@
(Shampoo’s estimation scheme based on Frobenius norm) The optimal solution of the Frobenius norm minimization $\min_{{\bm{S}}_{a},{\bm{S}}_{b}}\mathrm{Frob}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big):=\|\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]-{\bm{S}}\|_{\text{Frob}}$ with a two-sided precontioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}$ should satisfy the following condition.
@@
Remark: Although the solution can be obtained via rank-1 singular value decomposition (SVD) (Van Loan & Pitsianis, 1993) on this outer product, $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$, it can be computationally expensive to compute the solution due to the high dimensionality of the product. Moreover, the optimal solution is only achievable when the expectation of the outer product is computed exactly. Obtaining the optimal solution using SVD is even more expensive in tensor-valued cases.
@@
$\quad=\mathrm{Tr}\big({\bm{S}}_{a}^{2}\otimes{\bm{S}}_{b}^{2}-2\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(${\bm{S}}_{k}$ is symmetric for $k\in\{a,b\}$)}}$
$\quad=\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a Kronecker product)}}$
$\quad=\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big(({\bm{g}}{\bm{g}}^{\top})({\bm{S}}_{a}\otimes{\bm{S}}_{b})\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(linearity of the expectation)}}$
$\quad=\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}^{\top}\mathrm{vec}({\bm{S}}_{a}{\bm{G}}{\bm{S}}_{b})\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a Kronecker product)}}$
$\quad=\mathrm{Tr}\big({\bm{S}}_{a}^{2}\big)\mathrm{Tr}\big({\bm{S}}_{b}^{2}\big)-2\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{S}}_{a}{\bm{G}}{\bm{S}}_{b}\big)\big]+\text{const.}\quad\text{{\small\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}(Property of a trace)}}$
@@ Appendix G Key Distinction between Shampoo with trace scaling and KL-Shampoo
(Shampoo’s estimation scheme with trace scaling) The optimal solution of the von Neumann (VN) divergence (quantum relative entropy) minimization $\min_{{\bm{S}}_{a},{\bm{S}}_{b}}\mathrm{VN}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}],{\bm{S}}\big):=\mathrm{Tr}({\bm{S}})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\mathrm{LogM}({\bm{S}})\big)+\text{const.}$ with a two-sided precontioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}$ should satisfy the following condition.
@@
${\bm{S}}_{a}^{*}$ $=\mathrm{Diag}\big(\mathbb{E}[{\bm{G}}{\bm{G}}^{\top}]\big)=\mathrm{Diag}\big(\mathbb{E}[\big({\bm{G}}^{\odot 2}\big)\mathbf{1}]\big)$
${\bm{S}}_{b}^{*}$ $=\mathrm{Diag}\left(\frac{\mathbb{E}[{\bm{G}}^{\top}{\bm{G}}]}{\mathrm{Tr}\big(\mathbb{E}[{\bm{G}}{\bm{G}}^{\top}]\big)}\right)=\frac{\mathrm{Diag}\big(\mathbb{E}[\mathbf{1}^{\top}{\bm{G}}^{\odot 2}]\big)}{\mathrm{Tr}\big(\mathbb{E}[\mathbf{1}^{\top}\big({\bm{G}}^{\odot 2}\big)\mathbf{1}]\big)}=\frac{\mathrm{Diag}\big(\mathbb{E}[\mathbf{1}^{\top}\big({\bm{G}}^{\odot 2}\big)]\big)}{\sqrt{\mathrm{Tr}\big(\mathbb{E}[\mathbf{1}^{\top}\big({\bm{G}}^{\odot 2}\big)\big)\mathrm{Tr}\big(\mathbb{E}[\big({\bm{G}}^{\odot 2}\big)\mathbf{1}]\big)}}$
@@
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\mathrm{LogM}({\bm{S}})\big)+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}+{\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}\big)+\mathrm{Tr}\big(\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\big({\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}{\bm{g}}^{\top}\big(\mathrm{LogM}({\bm{S}}_{a})\otimes{\bm{I}}_{b}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{g}}{\bm{g}}^{\top}\big({\bm{I}}_{a}\otimes\mathrm{LogM}({\bm{S}}_{b})\big)\big)\big]+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a}){\bm{G}}{\bm{I}}_{b}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{I}}_{a}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a}){\bm{G}}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$=\mathrm{Tr}({\bm{S}}_{a})\,\mathrm{Tr}({\bm{S}}_{b})-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}{\bm{G}}^{\top}\mathrm{LogM}({\bm{S}}_{a})\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}\mathrm{LogM}({\bm{S}}_{b})\big)\big]+\text{const.}$
$=\mathrm{Tr}(\mathrm{ExpM}({\bm{P}}_{a}))\,\mathrm{Tr}(\mathrm{ExpM}({\bm{P}}_{b}))-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}{\bm{G}}^{\top}{\bm{P}}_{a}\big)\big]-\mathbb{E}\big[\mathrm{Tr}\big({\bm{G}}^{\top}{\bm{G}}{\bm{P}}_{b}\big)\big]+\text{const.}$ (15)